40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 9 23:12:56 2012, Jenne, Update, LSC, 11MHz reconnected to EOM 
    Reply  Wed Oct 10 02:15:14 2012, rana, Update, LSC, 11 MHz reconnected to EOM 
Message ID: 7513     Entry time: Tue Oct 9 23:12:56 2012     Reply to this: 7515
Author: Jenne 
Type: Update 
Category: LSC 
Subject: 11MHz reconnected to EOM 

Riju hasn't been in the lab in a long time to do any measurements, so I put the signals back to how they should be. 

I turned off / confirmed off the things which were sending signal to the EOM:  the network analyzer, the RF generator box, and the Marconi which supplies the 11MHz. 

I removed the cavity scanning cable, and terminated it, and put the regular 11MHz cable back on the splitter.

I then turned on the RF gen box and the Marconi.  The Marconi had been off, so we were not getting any 11MHz or 55MHz out of the RF gen. box.  This is why I couldn't lock any cavities last night (duh). 

On to locking!

----------------- In other news,

While swapping out the EOM cable, I noticed that the DC power supply sitting under the POX table was supplying a weird value, 17 point something volts.  I checked on the table to remind myself why that power supply is there...it's powering an RF amplifier right after the commercial PD that is acting as POP22.  The amplifier wants +15 and GND, so I reset the power supply to 15V.  We should add this to the list of things to fix, because it's dumb.  Either we need to put in the real POP22 (long term goal), or we need to get this guy some rack power, and do the same for any amplifiers for the Beat setup.  It's a little hoakey to have power supplies littering the lab.

ELOG V3.1.3-